# On the Verification of Control Flow Attestation Evidence

### v1

@@ 7.1 Control Flow Attestation and Auditing
The majority of existing work in this area focuses on the correctness and security of $\sf{\mathcal{P}rv}$ implementation to generate run-time evidence [4, 53, 54, 50, 63, 62, 51, 52, 49, 6, 58, 57]. They often mention that with authentic evidence, $\sf{\mathcal{V}rf}$ can detect the validity of the reported path by utilizing data structures computed prior to run-time (e.g., CFG+SS [53, 54, 50, 63, 6, 49] or hash-sets [62, 52, 4, 58]) or by emulating execution [51, 59]. However, as demonstrated by this work, all techniques except for the one utilizing verbatim evidence cannot infer the precise execution path on $\sf{\mathcal{P}rv}$ from their reported evidence. As such, they are not suitable for conducting root cause analysis after detecting the exploit on $\sf{\mathcal{P}rv}$.
@@
In terms of $\sf{\mathcal{V}rf}$ point-of-view, only two existing studies [11, 12] propose alternative techniques for the $\sf{\mathcal{V}rf}$ to determine path validity under different scenarios. ZEKRA [11] considers a context in which $\sf{\mathcal{V}rf}$ might also be untrusted and aims to learn secrets about the $\sf{\mathcal{P}rv}$’s program through the run-time evidence. To address this issue, they propose a protocol that outsources the verification of evidence to a semi-dishonest worker. This semi-dishonest worker evaluates the reported path by passing it through an arithmetic circuit representing the CFG and generating a (zero-knowledge) proof that the verification result was evaluated on the most recent path and the correct CFG. Unlike SABRE, ZEKRA changes the system model of CFA by assuming $\sf{\mathcal{V}rf}$ might be untrusted. ZEKRA’s zero-knowledge procedures are only concerned with determining path validity, while SABRE aims to conduct further root cause analysis from the reported path.
RAGE [12] presents a technique for path verification that does not require constructing a complete CFG. RAGE trains a Graph Neural Network (GNN) on a set of $\sf{\mathcal{P}rv}$’s known program traces and validates run-time evidence using the trained GNN. Because of this, RAGE presents a method in which no knowledge of the code (source code or binary) is required during the run-time validation on $\sf{\mathcal{V}rf}$, unlike SABRE and other techniques. In contrast, the verification in SABRE and CFG-based techniques does not require prior execution contexts. Although SABRE requires constructing a partial CFG, we show that this is feasible for low-end embedded applications, as apparent in Fig. 10(a). More importantly, RAGE does not support a root cause analysis from the received evidence, which is the main goal of this work.
7.2 Symbolic Execution/Root Cause Analysis
The concept of symbolic execution was first introduced given the challenge of testing the reliability of programs [77]. Prior to its introduction, common techniques were to select a finite set of inputs to test the program’s correctness (i.e., fuzzing) or perform a formal analysis of a program to determine if it meets a particular specification for all inputs. A challenge for these approaches comes from determining the set of inputs or formal specifications for evaluating the program. Symbolic execution provides an alternative approach by executing the program on symbolic inputs and evaluating the input conditions that lead to a particular execution path. As a result, one symbolic execution can represent an infinite set of executions on concrete inputs. Since this original work [77], further research addresses its challenges and demonstrates its use to verify security features/detect bugs.
Various tools exist to provide a symbolic execution engine [78, 79, 80, 81], and their ability to symbolically explore program states has been leveraged in prior works across multiple platforms to evaluate their security. Some examples include bug testing/vulnerability detection [82, 83, 84, 85, 86, 87, 88], root cause analysis [14, 15], and verification of security properties [89, 90]. Since symbolic execution can be complex due to challenges with modeling memory, maintaining states of complex (or infinite) programs, or modeling behaviors of inputs/peripherals, Concolic Execution [91, 92, 93] combines characteristics of program emulation and symbolic execution by leveraging concrete values to guide symbolic path exploration.
Root Cause Analysis [94, 95, 14, 96, 97, 98] is an automated process for identifying the source of software bugs/unexpected behavior in software. ARCUS [14] is a root cause analysis framework with the goal of identifying the sources of run-time exploits in production systems. ARCUS contains separate modules to detect buffer overflows (stack & heap), integer over/underflows, use-after-free vulnerabilities, and string format vulnerabilities. It uses a hardware tracing module, Intel PT, to record a trace of control flow transfers whenever a program executes. The root cause analysis starts when the underlying CFI technique notifies ARCUS of a control flow violation. ARCUS then collects the execution trace from Intel PT and uses it to emulate the malicious execution. To ensure accurate emulation, a snapshot of the initial memory state before the malicious execution is required, and this snapshot is taken whenever a new thread to run an application program is created. While sharing the same goal of the root cause identification, SABRE operates on much more limited information (i.e., without a snapshot of initial memory as it is not available from $\mathcal{C}F_{Slice}$) as well as targets different classes of devices (i.e., Intel desktops in ARCUS vs ARM/MSP430 embedded devices in SABRE).
@@ 7.2 Symbolic Execution/Root Cause Analysis
BENZENE [98] performs Root Cause Analysis starting from the application binary and crash-inducing inputs. It uses a data-flow graph (DFG) to track all values that were influenced by the inputs and mutates states along these paths to track when the program terminated successfully or crashed again. It then analyzes the crash and non-crash conditions to infer about crash conditions and generates a list of likely data-flow events that are the root causes. Unlike BENZENE, SABRE does not require data inputs and does not require constructing a DFG. Instead, SABRE combines CFG traversal with symbolic data-flow analysis to determine memory instructions that corrupt control data. Both SABRE and BENZENE create a report of a root cause. Yet, unlike BENZENE, SABRE generates and verifies a patch for the exploited root cause. Additionally, BENZENE assumes that data inputs were known, while SABRE relies on symbolic data-flow analysis to account for lacking data inputs.
8 Conclusion
@@
We study if and how run-time evidence produced by CFA architectures can be used to identify corrupted branch targets, pinpoint exploited memory instructions, and generate/verify patches. We analyze and classify types of run-time evidence generated by current CFA architectures and conclude that verbatim evidence provides $\sf{\mathcal{V}rf}$ with the most flexibility for remote analysis. We propose SABRE to verify the path validity and to enable $\sf{\mathcal{V}rf}$ to locate/patch vulnerabilities using only the application binary and run-time evidence. We implement and evaluate the public prototype of SABRE [16] to demonstrate its effectiveness and compatibility with hardware-based and TEE-based CFA architectures. Our results indicate the feasibility of exploit detection and root cause analysis from CFA-generated evidence (when reported in a verbatim/lossless manner), even in cases where source code is unavailable. Importantly, the type of remote analysis facilitated by CFA evidence would be infeasible locally at $\sf{\mathcal{P}rv}$. We believe that this initial exploration and results further encourage (much needed) research on the verification side of run-time attestation protocols. Among possible paths for future work, we highlight (1) root cause analysis and patching assuming vulnerability types other than buffer overflows or use-after-free-s (e.g., integer overflows/errors, double free); and (2) the detection and remediation of stealthier out-of-order execution attacks such as data-oriented programming.
Appendix A Control Flow Hijacking Examples
Listings 1 and 2 show examples in which $\sf{\mathcal{A}dv}$ can corrupt the target of an indirect branch without modifying/injecting code by exploiting a stack-based buffer overflow (Listing 1) and a use-after-free vulnerability (Listing 2).
A.1 Control Flow Attack via Buffer Overflow
Consider the stack overflow example in Listing 1 from CWE-121 [99] in which the buffer hostname is overflowed by $\sf{\mathcal{A}dv}$-controlled input. In this example, only 64 bytes are allocated on the stack for the buffer hostname. Yet, the $\sf{\mathcal{A}dv}$-controlled input user_supplied_addr points to the hostname that will be copied into that buffer. Therefore, $\sf{\mathcal{A}dv}$ could point to a hostname larger than 64 bytes. As a result, the $\sf{\mathcal{A}dv}$-controlled values would overflow into other stack variables when the hostname is copied into the buffer at line 10. Since hostname is allocated onto the stack, $\sf{\mathcal{A}dv}$ can launch a control flow attack by corrupting other indirect branch targets (e.g., return addresses or function pointers) that are also stored on the stack.
A.2 Control Flow Attack via Use-After-Free
Consider the example in Listing 2 inspired by Wickman et al. [100]. In this example, the function handle_net_input contains a use-after-free vulnerability. This function first allocates a new instance of the Parser struct (parser) onto the heap, which contains a function pointer handler (as seen in lines 6-10). After setting the fields in parser, a command cmd is read from the network on line 21. When cmd is INVALID, it executes a free of the parser object on line 24 and reads a new command from the network on line 25. If this next command is PARSE, a new Array object array is allocated on line 32, which is the same size as a Parser object. Consequently, array->buf now overlaps with the previous parser->handler function pointer. On line 34, input is received from the network to set array->buf. After that, the freed pointer parser->handler is used on line 36. As a result, an arbitrary memory address, based on the value of array->buf, is called. When $\sf{\mathcal{A}dv}$ controls the input used to set array->buf, they can arbitrarily decide which section of the code to execute due to the use-after-free of parser->handler.
1void host_lookup(char *user_supplied_addr){
2 struct hostent *hp;
3 in_addr_t *addr;
4 char hostname[64];
5 in_addr_t inet_addr(const char *cp);
6
7 validate_addr_form(user_supplied_addr);
8 addr = inet_addr(user_supplied_addr);
9 hp = gethostbyaddr(addr, sizeof(struct in_addr), AF_INET);
10 strcpy(hostname, hp->h_name);
11}
Listing 1: Stack-based Buffer Overflow Example [99]
1typedef struct{
2 long used;
3 char buf[24];
4} Array; // 32-byte
5
6typedef struct { // 32-byte
7 long status;
8 void *start, *current;
9 int (*handler)(void *buf);
10} Parser;
11
12enum Command { INVALID, PARSE, ... };
13int net_parser(void *buf);
14
15int handle_net_input(int client_fd) {
16 // allocation
17 Parser *parser = (Parser *)malloc(32);
18 parser->handler = &net_parser;
19
20 enum Command cmd = INVALID;
21 read(client_fd, &cmd, sizeof(cmd));
22
23 if (cmd == INVALID) {
24 free(parser);
25 read(client_fd, &cmd, sizeof(cmd));
26 }
27 .
28 .
29 .
30 if (cmd == PARSE) {
31 // re-allocation
32 Array *array = (Array *)malloc(32);
33 // parser->handler() changes
34 read(client_fd, array->buf, 24);
35 // use-after-free
36 parser->handler(array->buf);
37 free(parser);
38 }
39}
Listing 2: Use-after-free for Control Flow Hijacking

### v2

@@ 6. Related Work
The majority of existing work in this area focuses on the correctness and security of $\sf{\mathcal{P}rv}$ implementation to generate run-time evidence (Abera et al., 2016; Toffalini et al., 2019; Zhang et al., 2021; Sun et al., 2020; Wang et al., 2023; Dessouky et al., 2017, 2018; Zeitouni et al., 2017; Nunes et al., 2021b; Caulfield et al., 2023; Yadav and Ganapathy, 2023; Neto and Nunes, 2023; Armanuzzaman et al., 2025). They often mention that with authentic evidence, $\sf{\mathcal{V}rf}$ can detect the validity of the reported path by utilizing data structures computed prior to run-time (e.g., CFG+SS (Toffalini et al., 2019; Zhang et al., 2021; Sun et al., 2020; Wang et al., 2023; Caulfield et al., 2023; Nunes et al., 2021b) or hash-sets (Dessouky et al., 2017; Zeitouni et al., 2017; Abera et al., 2016; Yadav and Ganapathy, 2023)) or by emulating execution (Dessouky et al., 2018; Nunes et al., 2021a). As demonstrated by this work, all techniques that create hash-based/hybrid evidence (E1 and E3 from Sec. 3) cannot infer the precise malicious execution path on $\sf{\mathcal{P}rv}$ and are not suitable for conducting root cause analysis.
@@
Only two existing studies (Debes et al., 2023; Chilese et al., 2024) propose alternative approaches to CFA verification. ZEKRA (Debes et al., 2023) considers a context in which $\sf{\mathcal{V}rf}$ might also be untrusted and aims to learn secrets about $\sf{\mathcal{P}rv}$’s program through $\mathcal{C}F_{Log}$. To address this issue, ZEKRA outsources verification to a semi-dishonest worker, who uses an arithmetic circuit of CFG to generate a zero-knowledge proof of the verification process. ZEKRA’s procedures are only concerned with determining path validity, while SABRE aims to conduct further root cause analysis from $\mathcal{C}F_{Log}$. RAGE (Chilese et al., 2024) presents a technique for path verification that does not require constructing a complete CFG. RAGE trains a Graph Neural Network (GNN) on a set of $\sf{\mathcal{P}rv}$’s known $\mathcal{C}F_{Log}$-s and validates new $\mathcal{C}F_{Log}$-s using the trained GNN. Although SABRE requires constructing a partial CFG, we show this is feasible for low-end embedded applications, as apparent in Fig. 10(a). More importantly, RAGE does not support a root cause analysis from the received evidence, which is SABRE’s main goal.
Root Cause Analysis. Root-Cause Analysis (Zeller and Hildebrandt, 2002; Li et al., 2023; Yagemann et al., 2021b; Qasem et al., 2021; Liu et al., 2022; Park et al., 2024; Blazytko et al., 2020; Zeng et al., 2022; Zhou et al., 2022; Yagemann et al., 2021a) is an automated process for identifying the source of software bugs/unexpected behavior in software. Two closely related Root Cause Analysis frameworks are ARCUS and BENZENE.
ARCUS (Yagemann et al., 2021b) uses Intel PT to record an execution trace at run-time while a CFI scheme monitors execution until a control flow violation occurs. Then, it collects the execution trace and initial memory snapshot and sends them to a remote server to start emulating the malicious execution. During emulation, ARCUS performs checks for the root cause vulnerability. ARCUS and similar methods (Yagemann et al., 2021a; Zeng et al., 2022; Liu et al., 2022; Zhou et al., 2022) require a snapshot of the initial memory state. While sharing a similar goal, these prior approaches are not directly compatible with CFA. In CFA, $\sf{\mathcal{P}rv}$’s RoT may be unable to provide complete execution context, including initial memory snapshot and traces of all executing tasks since boot, particularly in on-demand attestation (Caulfield et al., 2023, 2024a; Neto and Nunes, 2023) in which $\sf{\mathcal{P}rv}$ either remains idle or executes unattested tasks in between CFA requests from $\sf{\mathcal{V}rf}$. Furthermore, storing a memory snapshot while the attested tasks execute would be infeasible for a resource-constrained $\sf{\mathcal{P}rv}$ MCU with limited memory (a common $\sf{\mathcal{P}rv}$ platform in CFA (Ammar et al., 2024)). To provide root cause analysis suitable with CFA evidence, SABRE operates without a snapshot of initial memory (which is not available from $\mathcal{C}F_{Slice}$).
@@
BENZENE (Park et al., 2024) performs its analysis using the application binary and crash-inducing inputs. It uses a data-flow graph (DFG) to track all influenced values and mutate states to track when the program terminated successfully or crashed again. Similarly to SABRE, BENZENE starts from the crash site and performs a backtracking process. It then analyzes the crash/non-crash conditions and generates a list of likely root cause data-flow events. Unlike BENZENE, SABRE does not require data inputs and does not require constructing a DFG. Instead, SABRE combines CFG traversal with symbolic data-flow analysis to determine exploited memory instructions without requiring data inputs. Unlike BENZENE, SABRE generates and verifies a binary patch for the identified root cause.
7. Conclusion
@@
We study if and how run-time evidence of CFA architectures can be used to identify corrupted branch targets, pinpoint exploited memory instructions, and generate/verify binary patches. We analyze and classify types of run-time evidence generated by current CFA architectures and conclude that verbatim evidence provides $\sf{\mathcal{V}rf}$ with the most flexibility for remote analysis. We introduce SABRE to verify path validity and to enable $\sf{\mathcal{V}rf}$ to locate/patch vulnerabilities using only the application binary and run-time evidence. We implement and evaluate the public prototype of SABRE (Caulfield et al., 2024b) to demonstrate its effectiveness and compatibility with hardware-based and TEE-based CFA architectures.